Kindly answer this question!!! Please!!! A car manufacturer offers 4 choices of body design, 10 color combination, and 3 different engines. How many different cars are required for a complete display?
Answer by Theo(13342)

You have 4 body designs, 10 color combinations, 3 engines.
For each body design you have 10 colors, so the total possible combinations are 4 * 10 = 40
For each one of these 40 combinations you have 3 engines, so the total possible combinations are 40 * 3 = 120.
-----
Your equation is:
4 * 10 * 3 = 120
-----
If I tell you that you have 2 body designs to choose from and 3 color combinations and 2 engines to choose from, you would have a total of 2 * 3 * 2 possible combinations. This can be shown as follows:
Let the 2 body designs be B1 and B2
Let the 2 colors be C1 and C2
Let the 2 engines be E1 and E2
-----
Your possible combinations are:
-----
B1C1E1
B1C1E2
B1C2E1
B1C2E2
B1C3E1
B1C3E2
B2C1E1
B2C1E2
B2C2E1
B2C2E2
B2C3E1
B2C3E2
-----
Your answer is 120 which is 4 * 10 * 3
-----
4 body designs
10 colors
3 engine types
-----